Problem 2: Let X be a binomially distributed random variable based on n 10 trials with success probability p 0.3. a) Compute P(X 3 8), P(x-7 and PX> 6) by hand, showing your work.
Answer #1

{\displaystyle f(k,n,p)=\Pr(k;n,p)=\Pr(X=k)={\binom {n}{k}}p^{k}(1-p)^{n-k}}

P(X <= 8) = 1 - P(X > 8) = 1 - ( P(X = 9) + P(X = 10)

1 - (10C9 0.3^9 * 0.7 + 0.3^10)

= 1 - (10 * 0.3^9 * 0.7 + 0.3^10)

= 0.999856

b)

P(X = 7)

= 10C7 * 0.3^7 * 0.7^3

= 0.009001

c)

P(X > 6)

= P(X = 7) + P(X = 8) + P(X = 9) + P(X = 10)

= 0.009001 + 10C9 0.3^8 *0.7^2 + (10 * 0.3^9 * 0.7 + 0.3^10)

= 0.010592